OpenStudy (anonymous):

factor x^3-1 is this a prime or if not how would I factor

OpenStudy (anonymous):

This is the difference of 2 perfect cubes which can be factored. a^3 - b^3 = (a - b) (a^2 + ab + b^2). Here, a=x, and b=1.